Hargrave Ltd has just paid a dividend of $2.00 per share on its stock. The dividends are expected to grow at a constant 9% p.a. indefinitely. Calculate the current price if investors require a 14% p.a. return on Hargrave stock. What will the price be in 3 years? In 15 years?
Answer #1

Using the Gordon Model

Situation 1 - Dividend is expected to grow @9% indefinitely

Price (Time 0) = Dividend(time 1)/Expected Rate of Return (ke) - Growth Rate(g)

=2(1.09)/0.14-0.09

=43.6

Situation 2 - Price in 3 years

Price (Time 3) = Dividend(time 4)/ ke - g

=2(1.09)^4/0.14-0.09

=2.8232/0.05

=56.46

Situation 3 - Price in 15 years

Price (Time 15) = Dividend (time 16)/ ke - g

=2(1.09)^16/0.14-0.09

=7.9406/0.05

=158.8122